# supplement

### v1

@@ 5 Conclusions
The main contribution of this paper is an optimization-based bound for kernel-based estimates that is tight, even in the non-asymptotic, low-data regime, and that can handle correlated noise sequences. By modeling both the latent function and the noise as deterministic functions in the RKHS framework, it leverages knowledge on their respective maximum RKHS norm. For the latent function, this is a common assumption in the literature on kernel-based uncertainty bounds; for the noise function, we discuss how it generalizes the common setting of energy-bounded noise. Still, modeling the noise as a deterministic quantity complicates leveraging multiple measurements for the same input locations; overcoming Assumption 2 will be subject of future research.
The obtained bounds suffer from the common criticalities and limitations of kernel-based learning, which are (a) dealing with kernel mis-specification; (b) knowing the correct value of the RKHS-norm bound; (c) handling large data-sets. Both issues (a) and (b) are typically addressed empirically, the first by hyper-parameter tuning via cross-validation [Wahba, 1990] and the second by estimating the bound value from data [Tokmak et al., 2024]. However, further research will be devoted in rigorously assessing the robustness of the obtained bounds with respect to possible mis-specifications in (a) and (b). Issue (c) is the most critical, as it can impact the computational time and the matrix conditioning, possibly deteriorating the bound quality. To address it, kernel approximations can be deployed, and their impact on the overall bound should be carefully studied. In regard to large data-sets, we would also like to emphasize that the obtained bound is not suitable for asymptotic analyses, but is of great practical utility in view of downstream tasks.
@@
This work was supported by the European Union’s Horizon 2020 research and innovation programme, Marie Skłodowska-Curie grant agreement No. 953348, ELO-X. AL thanks Manish Prajapat for helpful discussions.
@@ Appendix A Finite-dimensional representation of optimization problems
In this Section we first prove that optimization problems (2) and (4) admit a finite-dimensional representation (Lemma A.1 and Lemma A.2 in Section A.1). Next, in Section A.2 we present two coordinate transformations that will be deployed in the remaining sections.
@@ A.1 Representer Theorems
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}\alpha^{f}\in\mathbb{R}^{N+1},\\ c^{w}\in\mathbb{R}^{N},\end{subarray}}\quad$ $K^{f}_{N+1,1:N+1}\alpha^{f}$ (A.2a)
@@
Let $\mathbb{X}=\{x_{1},\ldots,x_{N}\}$ be the set of training input locations and $\mathbb{X}_{+}=\mathbb{X}\cup\{x_{N+1}\}$, the same set augmented with the test point. We denote by $\mathcal{H}_{k^{f}}^{\parallel}=\{f\in\mathcal{H}_{k^{f}}:f\in\mathrm{span}(k^{f}(\cdot,x_{i}),x_{i}\in\mathbb{X}_{+})\}$ the span of kernel functions evaluated at the training and test input locations, as well as by $\mathcal{H}_{k^{f}}^{\perp}$ its orthogonal complement, i.e., $\mathcal{H}_{k^{f}}^{\perp}=\{f^{\perp}\in\mathcal{H}_{k^{f}}:\langle f^{\perp},f^{\parallel}\rangle_{\mathcal{H}_{k^{f}}}=0\text{ for all }f^{\parallel}\in\mathcal{H}_{k^{f}}^{\parallel}\}$. Hence, any function $f\in\mathcal{H}_{k^{f}}$ can be written as $f=f^{\parallel}+f^{\perp}$, where $f^{\parallel}\in\mathcal{H}_{k^{f}}^{\parallel}$ and $f^{\perp}\in\mathcal{H}_{k^{f}}^{\perp}$. Note that the cost of the optimization problem is $f(x_{N+1})=\langle f,k(x_{N+1},\cdot)\rangle_{\mathcal{H}_{k^{f}}}$, which is insensitive to the orthogonal part $f^{\perp}$. Regarding the constraints, note that all functions $f^{\perp}\in\mathcal{H}_{k^{f}}^{\perp}$ do not affect the equality constraint (2b) while tightening the inequality constraints (2c); hence, it is optimal to set $f^{\perp}\equiv 0$. By the same arguments, it is optimal to set $w^{\perp}\equiv 0$. where the orthogonal complement is defined with respect to the finite-dimensional subspace $\mathcal{H}_{k^{w}}^{\parallel}=\{w\in\mathcal{H}_{k^{w}}:w\in\mathrm{span}(k^{w}(\cdot,x_{i}),x_{i}\in\mathbb{X})\}$, which excludes $k^{w}(\cdot,x_{N+1})$ as the cost is insensitive to $w(x_{N+1})$, the value of the noise function at the test point. Hence, it follows that for all functions $f\in\mathcal{H}_{k^{f}}$ and $w\in\mathcal{H}_{k^{w}}$, the respective orthogonal parts $f^{\perp}\in\mathcal{H}_{k^{f}}^{\perp}$ and $w^{\perp}\in\mathcal{H}_{k^{w}}^{\perp}$ can be set to zero without affecting feasibility or optimality of the candidate function.
@@
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}\alpha^{f}\in\mathbb{R}^{N+1},\\ c^{w}\in\mathbb{R}^{N},\end{subarray}}\quad$ $K^{f}_{N+1,1:N+1}\alpha^{f}$ (A.3a)
@@
Analogous to Lemma A.1, it holds that setting $f^{\perp}\equiv 0$ retains optimality of any candidate function $f\in\mathcal{H}_{k}$, with $f=f^{\parallel}+f^{\perp}$. Similarly for the noise, it holds that $w^{\perp}\equiv 0$. Attainment of the supremum is also established along the lines of Lemma A.1, noting that the sum of norm-constraints (4c) defines a closed and bounded set in $\mathcal{H}_{k^{f}}^{\parallel}\times\mathcal{H}_{k^{w}}^{\parallel}$. Finally, the finite-dimensional optimization problem follows from replacing $f,w$ with their finite-dimensional expressions (A.1). ∎
@@ A.2 Coordinate transformations
We now present two transformations that will allow us to simplify the finite-dimensional representations (A.2) and (A.3).
@@
with $\delta_{1}\in\mathbb{R}^{N}$, $\delta_{2}\in\mathbb{R}^{r}$, which allows compute the components of the solution determined by the training data. For clarity, we emphasize that the partitioning of the matrices in Eq. A.13 on the left- and right-hand side is different: the first line on the left-hand side contains $N$ rows, th first line on the right-hand side, $r$ rows.
Appendix B Proof of Lemma 1
Starting from the finite-dimensional formulation of the relaxed problem (4) as given in (A.3), we first apply the two coordinate transformations presented in Section A.2 (Section B.1). This allows us to obtain a simplified problem formulation — a linear program with a single norm-ball constraint — that can be solved directly (Section B.2). Finally, we also present the result for the lower bound (Section B.3).
@@ B.1 Preliminary coordinate transformation
=(A.10)${{=}}\begin{bmatrix}R^{\top}&0\end{bmatrix}\begin{bmatrix}Q_{11}^{\top}&Q_{21}^{\top}\\ Q_{12}^{\top}&Q_{22}^{\top}\end{bmatrix}\begin{bmatrix}I_{N}&0\\ 0&(\sigma R^{w}_{1:N,1:N})^{-1}\end{bmatrix}\begin{bmatrix}\theta\\ c_{w}\end{bmatrix}$
=(A.13)${{=}}R^{\top}\delta_{1}.$ (B.1)
@@ B.2 Analytical solution
$\Phi_{N+1}Q_{11}\delta_{1}^{\star,\sigma}$ =(B.7b)${=}(\Phi_{N+1}\Phi_{1:N}^{\top})R^{-1}R^{-\top}y,$
=(B.7c)${=}K^{f}_{N+1,1:N}\left(K^{f}_{1:N,1:N}+\sigma^{2}K^{w}_{1:N,1:N}\right)^{-1}y,$
=(5a)${=}f^{\mu}_{\sigma}(x_{N+1}).$ ) .
@@
$\|\delta_{1}^{\star,\sigma}\|_{2}^{2}$ =(B.1)${=}y^{\top}R^{-1}R^{-\top}y$
=(B.7c)${=}y^{\top}\left(K^{f}_{1:N,1:N}+\sigma^{2}K^{w}_{1:N,1:N}\right)^{-1}y$
=(6)${=}\|g^{\mu}_{\sigma}\|_{\mathcal{H}_{k^{f}+k^{\sigma}}}^{2}.$
@@ B.3 Optimal relaxed solution for the lower bound
Let Assumptions 1 and 2 be satisfied. Then, for all $\sigma\in(0,\infty)$, it holds that
@@
Appendix C Proof of Theorem 1
@@
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}\theta\in\mathbb{R}^{r},\end{subarray}}\quad$ $\Phi_{N+1}\theta$ (C.1a)
@@
We will analyze the solution of Problem (C.1) for different active sets. Here, the term “active set” refers to a subset of the RKHS-norm constraints (C.1b) and (C.1c) that are strictly active, i.e., influence the optimal primal solution of the problem. For strictly active constraints (C.1b) and (C.1c) there exist respective Lagrangian multipliers [^2] Note that if the Lagrangian multipliers for a given primal solution are non-unique, then it is possible that multiple cases apply concurrently; however, this does not affect our analysis., $\lambda^{f}$ and $\lambda^{w}$, that are strictly positive. We investigate the following combinations:
@@ 4. Case 4
Based on the solutions for fixed active sets, the optimal solution can then be found by case distinction. We discuss each case separately, obtaining the corresponding analytical solution, presenting the feasibility check and elucidating the connection with the solution of the relaxed problem in Sections C.1, C.2, C.3 and C.4; note that Sections C.2 and C.3 provide the proofs for Propositions 1, 2 and 1. We then show how to practically check which set is active, and obtain the desired claim (9) in Section C.5. We conclude the section by presenting the result for the lower bound (Section C.6).
@@ C.1 Case 1: Noise constraint inactive
We now consider the case in which only (C.1b) is strictly active, proving Proposition 1.
@@
For $\sigma\rightarrow\infty$, for the relaxed solution $\overline{f}^{\sigma}(x_{N+1})$ in Appendix B, it holds that
@@ C.2 Case 2: Function constraint inactive
We proceed by considering the case in which only (C.1b) is active, proving the result given in Proposition 2.
@@
=(C.6)${=}\left\|(R^{w}_{1:N,1:N})^{-1}y-\tilde{Q}\begin{bmatrix}\tilde{R}\\ 0\end{bmatrix}\theta\right\|_{2}^{2}$
=(C.7a)${=}\left\|\tilde{Q}^{\top}(R^{w}_{1:N,1:N})^{-1}y-\begin{bmatrix}\tilde{R}\\ 0\end{bmatrix}\theta\right\|_{2}^{2}$
=(C.7d)${=}\left\|\begin{bmatrix}\tilde{Q}_{11}^{\top}&\tilde{Q}_{21}^{\top}\end{bmatrix}(R^{w}_{1:N,1:N})^{-1}y-\tilde{R}\theta\right\|_{2}^{2}$
@@
$\mathrm{s.t.}\quad$ $\|z_{1}\|_{2}^{2}\leq\left(\Gamma_{w}^{2}-\tilde{y}^{\top}\left(K^{w}_{1:N,1:N}-M\right)\tilde{y}\right).$ (C.12b)
Noting that, by Assumption 1, the right-hand side of the constraint (C.12b) is non-negative, the optimal solution of the above problem is given as
@@
=(C.7c),(C.10)${{=}}\|P\Phi_{N+1}^{\top}\|_{P^{-1}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}+\Phi_{N+1}\theta^{\mu},$ , (C.15)
@@
=(C.7c)${{=}}\left(\Phi_{1:N}^{\top}(K^{w}_{1:N,1:N})^{-1}\Phi_{1:N}\right)^{-1}\Phi_{1:N}^{\top}\left(\tilde{y}+\lambda\frac{\sqrt{\Gamma_{w}^{2}-\tilde{y}^{\top}\left(K^{w}_{1:N,1:N}-M\right)\tilde{y}}}{\sqrt{\lambda^{\top}M\lambda}}\right)$ )
@@
=(5b)${{=}}\lim_{\sigma\rightarrow 0}\frac{1}{\sigma^{2}}\Sigma^{f}_{\sigma}(x_{N+1}).$ ) . (C.17)
@@
$=\lim_{\sigma\rightarrow 0}\sigma^{2}\left(\Gamma_{f}^{2}+\frac{\Gamma_{w}^{2}}{\sigma^{2}}-\frac{1}{\sigma^{2}}\tilde{y}^{\top}\left(K^{w}_{1:N,1:N}-M\right)\tilde{y}\right)$
@@
Proof of Corollary 1
We now prove Corollary 1, which simplifies the general result of Proposition 2 under the assumptions that the kernel matrix $K^{f}_{1:N,1:N}$ is invertible and the test point is equal to the $k$-th training point, i.e., $x_{N+1}=x_{k}$, for some $k\in\{1,\ldots,N\}$. In this case, the $k$-th and $(N+1)$-th row of the kernel matrix $K^{f}_{1:N+1,1:N+1}$ are identical. In terms of the singular value decomposition, by (A.5) this implies that
@@
i.e., the relation $\Phi_{N+1}^{\top}=\Phi_{1:N}^{\top}\lambda$ holds for $\lambda=e_{k}$, with $e_{k}$ being the $k$-th unit vector. Since $\Phi_{1:N}\in\mathbb{R}^{r\times N}$ has rank $r=N$, it is invertible. This allows to simplify the expression for the optimal cost using that $M\tilde{y}=y$, with $M$ and $\tilde{y}$ defined as in Eqs. C.9 and C.10:
@@ C.4 Case 4: Both constraints inactive
=$=\begin{cases}0,&\text{if $\Phi_{N+1}=0$,}\\ \infty,&\text{if $\Phi_{N+1}\neq 0$.}\end{cases}$
@@ C.5 Finding the correct active set
$\overline{f}(x_{N+1})=\begin{cases}\sqrt{K^{f}_{N+1,N+1}}\Gamma_{f},&\text{in Cases~{}1 and 4, }\\ \Phi_{N+1}\theta^{\mu}+\|P\Phi_{N+1}^{\top}\|_{P^{-1}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}&\text{in Cases 2 and 4, }\\ f^{\mu}_{\sigma^{\star}}(x_{N+1})+\sqrt{\Gamma_{f}^{2}+\frac{\Gamma_{w}^{2}}{(\sigma^{\star})^{2}}-\|g^{\mu}_{\sigma^{\star}}\|_{\mathcal{H}_{k^{f}+k^{\sigma^{\star}}}}^{2}}\sqrt{\Sigma^{f}_{\sigma^{\star}}(x_{N+1})},&\text{in Cases 3 and 4.}\end{cases}$
=$=\begin{cases}\lim_{\sigma\rightarrow\infty}\overline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|y-K^{f}_{1:N,N+1}\frac{\Gamma_{f}}{\sqrt{K^{f}_{N+1,N+1}}}\right\|_{(K^{w}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{w}^{2}$} },\\ \lim_{\sigma\rightarrow 0}\overline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|\theta^{\mu}+\frac{P\Phi_{N+1}^{\top}}{\|P\Phi_{N+1}^{\top}\|_{P^{-1}}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}\right\|_{2}^{2}\leq\Gamma_{f}^{2}$}, }\\ \inf_{\sigma\in(0,\infty)}\overline{f}^{\sigma}(x_{N+1}),&\text{otherwise.}\end{cases}$

### v3

@@ 5 Conclusions
The main contribution of this paper is an optimization-based, distribution-free bound for kernel-based estimates that is tight, even in the non-asymptotic, low-data regime, and that can handle correlated and biased noise sequences. The proposed bound generalizes known results from kernel interpolation in the noise-free setting and from linear regression under energy-bounded noise. In the case of bounded sub-Gaussian noise, the numerical results highlight the competitiveness of the bound with existing probabilistic bounds in terms of its conservatism, and showcase its high potential for safe control as a downstream task. Moreover, the experiments highlight how the deterministic nature of the proposed bound enables the rigorous certification of subset-of-data selection or mixture-of-experts strategies to handle large data sets. Future work may investigate the effectiveness of the proposed bound for further downstream tasks, such as Bayesian optimization or model-based reinforcement learning.
@@
This work was supported by the European Union’s Horizon 2020 research and innovation programme, Marie Skłodowska-Curie grant agreement No. 953348, ELO-X. The authors thank the anonymous reviewers for their constructive comments. AL thanks Philipp Hennig, Motonobu Kanagawa and Manish Prajapat for helpful discussions.
@@ Appendix A Finite-dimensional representation of optimization problems
In this Section we first prove that optimization problems (2) and (4) admit a finite-dimensional representation (Lemma˜A.1 and Lemma˜A.2 in Section˜A.1). Next, in Section˜A.2 we present two coordinate transformations that will be deployed in the remaining sections.
@@ A.1 Representer Theorems
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}\alpha^{f}\in\mathbb{R}^{N+1},\\ c^{w}\in\mathbb{R}^{N}\end{subarray}}\quad$ $K^{f}_{N+1,1:N+1}\alpha^{f}$ (A.2a)
@@
[…]
@@
$\overline{f}^{\sigma}(x_{N+1})=\sup_{\begin{subarray}{c}\alpha^{f}\in\mathbb{R}^{N+1},\\ c^{w}\in\mathbb{R}^{N}\end{subarray}}\quad$ $K^{f}_{N+1,1:N+1}\alpha^{f}$ (A.3a)
@@
Analogous to Lemma˜A.1, it holds that setting $f^{\perp}\equiv 0$ retains optimality of any candidate function $f\in\mathcal{H}_{k}$, with $f=f^{\parallel}+f^{\perp}$. Similarly for the noise, it holds that $w^{\perp}\equiv 0$. Attainment of the supremum is also established along the lines of Lemma˜A.1, noting that the sum of norm-constraints (4c) defines a closed and bounded set in $\mathcal{H}_{k^{f}}^{\parallel}\times\mathcal{H}_{k^{w}}^{\parallel}$. Finally, the finite-dimensional optimization problem follows from replacing $f,w$ with their finite-dimensional expressions (A.1). ∎
@@ A.2 Coordinate transformations
We now present two transformations that will allow us to simplify the finite-dimensional representations (A.2) and (A.3). The first one will be used to deal with the possible rank-deficiency of the kernel matrix, the second one, to decompose the hypothesis space into orthogonal features. A subset of the corresponding weights will be fully determined by the training data, while the remaining ones will be adversarially chosen to obtain the worst-case value of the latent function at the test point. We point the interested reader to Müller and Schaback [2009]; Pazouki and Schaback [2011] for details on similar basis transformations for kernel spaces.
@@
with $\delta_{1}\in\mathbb{R}^{N}$, $\delta_{2}\in\mathbb{R}^{r}$, which allows compute the components of the solution determined by the training data. For clarity, we emphasize that the partitioning of the matrices in Eq.˜A.13 on the left- and right-hand side is different: the first line on the left-hand side contains $N$ rows, the first line on the right-hand side, $r$ rows.
Appendix B Proof of Lemma˜1
Starting from the finite-dimensional formulation of the relaxed problem (4) as given in (A.3), we first apply the two coordinate transformations presented in Section˜A.2 (Section˜B.1). This allows us to obtain a simplified problem formulation — a linear program with a single norm-ball constraint — that can be solved directly (Section˜B.2). Finally, we also present the result for the lower bound (Section˜B.3).
@@ B.1 Preliminary coordinate transformation
[…]
@@ B.2 Analytical solution
[…]
@@
[…]
@@ B.3 Optimal relaxed solution for the lower bound
Let Assumptions˜1 and 2 be satisfied. Then, for all $\sigma\in(0,\infty)$, it holds that
@@
Appendix C Proof of Theorem˜1
@@
$\overline{f}(x_{N+1})=\sup_{\begin{subarray}{c}\theta\in\mathbb{R}^{r}\end{subarray}}\quad$ $\Phi_{N+1}\theta$ (C.1a)
@@
We will analyze the solution of Problem (C.1) for different active sets. Here, the term “active set” refers to a subset of the RKHS-norm constraints (C.1b) and (C.1c) that are strictly active, i.e., influence the optimal primal solution of the problem. For strictly active constraints (C.1b) and (C.1c) there exist respective Lagrangian multipliers, $\lambda^{f}$ and $\lambda^{w}$, that are strictly positive. We investigate the following combinations:
@@ 4. Case 4
Based on the solutions for fixed active sets [^3] The presented analysis of Cases 1-3 considers a slightly more permissive setting, which allows some of the Lagrange multipliers to be zero, i.e., the corresponding constraint to be inactive or weakly active. Thus, in some scenarios multiple cases might be applicable (see Section C.5); yet, this does not affect our analysis as the cases cover all possible scenarios. , the optimal solution can then be found by case distinction. We discuss each case separately, obtaining the corresponding analytical solution, presenting the feasibility check and elucidating the connection with the solution of the relaxed problem in Sections˜C.1, C.2, C.3 and C.4; note that Sections˜C.2 and C.3 provide the proofs for Propositions˜1, 2 and 1. We then show how to practically check which set is active, and obtain the desired claim (9) in Section˜C.5. We conclude the section by presenting the result for the lower bound (Section˜C.6).
@@ C.1 Case 1: Noise constraint inactive
We now consider the case in which only (C.1b) is active, proving Proposition˜1.
@@
For $\sigma\rightarrow\infty$, for the relaxed solution $\overline{f}^{\sigma}(x_{N+1})$ in Appendix˜B, it holds that
@@ C.2 Case 2: Function constraint inactive
We proceed by considering the case in which only (C.1b) is active, proving the result given in Proposition˜2.
@@
[…]
@@
[…]
Noting that, by Assumption˜1, the right-hand side of the constraint (C.12b) is non-negative, the optimal solution of the above problem is given as
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Proof of Corollary˜1
We now prove Corollary˜1, which simplifies the general result of Proposition˜2 under the assumptions that the kernel matrix $K^{f}_{1:N,1:N}$ is invertible and the test point is equal to the $k$-th training point, i.e., $x_{N+1}=x_{k}$, for some $k\in\{1,\ldots,N\}$. In this case, the $k$-th and $(N+1)$-th row of the kernel matrix $K^{f}_{1:N+1,1:N+1}$ are identical. In terms of the singular value decomposition, by (A.5) this implies that
@@
i.e., the relation $\Phi_{N+1}^{\top}=\Phi_{1:N}^{\top}\lambda$ holds for $\lambda=e_{k}$, with $e_{k}$ being the $k$-th unit vector. Since $\Phi_{1:N}\in\mathbb{R}^{r\times N}$ has rank $r=N$, it is invertible. This allows to simplify the expression for the optimal cost using that $M\tilde{y}=y$, with $M$ and $\tilde{y}$ defined as in Eqs.˜C.9 and C.10:
@@
Remark.
Corollary˜1 has been derived as a particular case of Proposition˜2, which provides the analytic solution for the case $\sigma\to 0$ occurring when the kernel matrix is rank-deficient. From this perspective, the scenario in which the test-input belongs to the training data-set is one of the particular situations leading to a drop in the rank of the kernel matrix. However, the proof of Corollary˜1 could be alternatively carried out following the steps of the one of Proposition˜1.
@@ C.4 Case 4: Both constraints inactive
$=\begin{cases}0,&\text{if $\Phi_{N+1}=0$,}\\ \infty,&\text{if $\Phi_{N+1}\neq 0$.}\end{cases}$
@@ C.5 Finding the correct active set
[…]
$=\begin{cases}\lim_{\sigma\rightarrow\infty}\overline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|y-K^{f}_{1:N,N+1}\frac{\Gamma_{f}}{\sqrt{K^{f}_{N+1,N+1}}}\right\|_{(K^{w}_{1:N,1:N})^{-1}}^{2}\leq\Gamma_{w}^{2}$} },\\ \lim_{\sigma\rightarrow 0}\overline{f}^{\sigma}(x_{N+1}),&\text{if \mbox{$\left\|\theta^{\mu}+\frac{P\Phi_{N+1}^{\top}}{\|P\Phi_{N+1}^{\top}\|_{P^{-1}}}\sqrt{\Gamma_{w}^{2}-y^{\top}\left(K^{w}_{1:N,1:N}\right)^{-1}y+\|\theta^{\mu}\|_{P^{-1}}^{2}}\right\|_{2}^{2}\leq\Gamma_{f}^{2}$}, }\\ \inf_{\sigma\in(0,\infty)}\overline{f}^{\sigma}(x_{N+1}),&\text{otherwise.}\end{cases}$